# The Bayesian Geometry of Transformer Attention

### v2

@@ 4.3. Length Generalization Requires Late-Layer Attention
To identify which components support stable rollout, we train a variant transformer in which attention is disabled in the top two layers but FFNs and residuals remain intact.
The no-late-attention model fits the training horizon reasonably well ($1.57\times 10^{-3}$ bits), but breaks down under rollout:
@@
4.4. MLP Controls: Architectural Necessity of Attention
Capacity-matched MLPs trained under identical conditions fail in both wind tunnels.
Bijections.
The MLP achieves MAE $\approx 1.85$ bits—-about $618\times$ worse than the transformer—- and shows no improvement from 100k to 150k steps. Its entropy curve remains nearly flat, indicating it learns only the marginal output distribution.
HMMs.
The MLP achieves MAE $\approx 0.40$ bits at all lengths (Table 1), showing no sign of recursive computation. The flat per-position error profile (Figure 10) indicates collapse to a position-averaged approximation rather than belief tracking.
These failures cannot be attributed to optimization, data, or capacity. They reflect the absence of content-addressable routing and residual compositionality—- key geometric ingredients supplied by attention.
@@ 4.5. Multi-Seed Consistency
To ensure that Bayesian tracking is not an artifact of initialization or optimization noise, we repeated all HMM experiments across five independent random seeds. Per-position error curves for all seeds (Figure 11) nearly overlap at $K=20$, $K=30$, and $K=50$.
The seed-to-seed variability is negligible compared to the gap between transformer and MLP performance, confirming that the learned Bayesian algorithm is robust to initialization and training noise.
Figure 1. Bijection wind tunnel: transformer matches the Bayesian posterior; MLP cannot. Entropy trajectories at 150k training steps. The transformer lies essentially on top of the analytic Bayes curve across positions, while the capacity-matched MLP barely reduces uncertainty and fails to implement hypothesis elimination. This is the comparison summarized quantitatively in Table 1 and discussed in Section 4.1.
Figure 2. Bijection wind tunnel: per-sequence entropy dynamics. Eight randomly chosen bijections from the test set. Each panel shows transformer entropy (solid) and analytic Bayes entropy (dashed) as a function of position. The sawtooth pattern—-discrete drops when mappings are revealed and collapses to (near) zero when previously seen inputs reappear—-confirms that the transformer is performing stepwise hypothesis elimination, not merely matching the Bayes curve in aggregate.
@@
Figure 11. Multi-seed robustness of HMM length generalization. Overlay of per-position transformer MAE curves across five random seeds for $K=20$, $K=30$, and $K=50$. Seed-to-seed variability is negligible relative to the transformer—-MLP gap, showing that the learned Bayesian algorithm is robust to initialization and optimization noise.
@@
Figure 16. Value-manifold unfurling during training. PCA projection of attention outputs in the bijection model, colored by analytic posterior entropy. At 100k steps, low-entropy states are tightly clustered; by 150k, they lie along a smooth one-dimensional curve parameterized by entropy, enabling fine-grained encoding of posterior states.Each point is an attention output (head output or block attention output —- whichever you used) at a supervised prediction position; PCA is fit on the pooled outputs and then plotted, colored by analytic posterior entropy.
@@ 5.1. Layer 0 Creates the Hypothesis Frame
The computation begins with a structural operation: Layer 0 attention constructs the hypothesis space in which all subsequent inference takes place. Keys at this layer form an approximately orthogonal basis over input tokens (Figure 14), providing a coordinate system over which posterior mass can be represented and manipulated.
Head-wise ablations confirm the indispensability of this step. A single Layer 0 “hypothesis-frame head” dominates the layer’s contribution (Figure 4), and ablating this head alone severely disrupts calibration. Here “hypothesis-frame head– means the head whose keys span the near-orthogonal basis over hypothesis tokens and whose values instantiate the corresponding per-hypothesis slots in the residual stream. No other attention head exhibits comparable sensitivity. This identifies a structural bottleneck: forming the hypothesis frame is a prerequisite for any later Bayesian computation.
Once established, this frame remains stable through training. Attention maps at Layer 0 change little across checkpoints, even as the value manifold and calibration improve substantially. The model therefore learns the geometry of the inference problem early, and subsequently refines numerical precision within this fixed frame.
@@ 5.2. Sequential Bayesian Elimination Across Depth
Layer-wise ablations (Figure 3) show that removing any single layer (attention + FFN, as implemented) increases calibration error by more than an order of magnitude. This demonstrates that the computation is not shallow or redundant. Each layer provides a distinct and non-interchangeable refinement step, forming a sequential, compositional realization of Bayesian updates.
Together, these observations indicate that transformers implement Bayesian elimination not via a single transformation, but through a depth-wise sequence of projections and refinements within the Layer 0 frame.
@@ 5.3. Attention as Content-Addressable Routing
Across all depths, attention serves a consistent geometric role: it retrieves the components of the belief state relevant for the next update.
Three observations support this routing interpretation:
@@
  • Stable routing during late refinement (Figures 16 and 17) shows that once the frame is correct, attention maps change minimally even as calibration improves.
@@ 5.4. Value-Space Manifolds and Precision Refinement
After routing stabilizes, the final layers refine the precision of the posterior representation. Figures 16 and 17 show that:
@@
This geometric unfurling enables fine-grained encoding of posterior confidence and accounts for late-position improvements in calibration. Importantly, this refinement occurs while attention maps remain nearly unchanged, producing a clear frame–precision dissociation: attention defines where information flows, while downstream transformations refine how precisely beliefs are encoded.
@@ 6.2. Depth as Compositional Necessity
A central conclusion from the ablation studies is that depth is not redundant. In both wind tunnels, removing any individual layer increases calibration error by more than an order of magnitude (Figure 3). This shows that Bayesian reasoning is expressed as a sequence of compositional projections, each layer refining the belief state in a way that cannot be collapsed into a single transformation.
This stands in contrast to wide, shallow architectures: even with comparable parameter counts and identical training, MLPs fail to perform hypothesis elimination or state tracking (Section 4.4). Bayesian inference requires hierarchical refinement, and transformers supply the appropriate inductive bias through depth and residual composition.
@@ 6.3. From Wind Tunnels to Natural Language
While the wind tunnels are deliberately simplified, they capture the essential structure of probabilistic inference: integrating evidence over time to update latent beliefs. Large language models operate in a far more complex setting, with high-dimensional latent spaces and ambiguous, multi-modal evidence. Yet the geometric ingredients observed here—-orthogonal hypothesis axes, depth-wise refinement, and stable routing—-are structural rather than task-specific.
The results therefore suggest that the probabilistic behaviors exhibited by LLMs may arise not only from scale or data richness but also from architectural geometry. Wind tunnels provide a verifiable lower bound: they show that transformers can implement Bayesian inference exactly when the posterior is known.
6.4. Architectural Necessity and MLP Failure
The capacity-matched MLP controls clarify which architectural components are essential. Even with similar parameter counts and identical data exposure, MLPs fail catastrophically in both wind tunnels, with entropy errors on the order of $0.4$ bits (Table 1). These failures are not due to optimization difficulties: the tasks are simple, gradients are well-behaved, and training converges smoothly.
Instead, the gap reflects the absence of:
  • content-addressable retrieval of hypotheses,
  • compositional refinement through depth, and
  • stable routing structures that support long-horizon inference.
Transformers succeed because attention supplies the geometric mechanisms—-orthogonal bases, selective routing, and progressive focusing—-that MLPs lack. The failure of matched MLPs therefore serves as a clean demonstration that attention is essential for Bayesian structure learning in context, at least among the architectures tested.
@@ 6.5. A Lower Bound for Reasoning in LLMs
The wind tunnels establish a principled baseline for mechanistic reasoning in transformers. If a model cannot implement Bayes in a setting with a closed-form posterior and impossible memorization, it offers little evidence of genuine inference capability in natural language. Conversely, the fact that small, verifiable transformers succeed here—-with interpretable geometric mechanisms—-suggests that similar structures may underpin reasoning in large models.
This provides a concrete research direction: search for the same geometric signatures in frontier LLMs. The diagnostics used here—-key orthogonality, QK sharpening, value-manifold structure, and routing stability—-offer testable predictions for analyzing pretrained language models.
@@ 7.1. Bayesian Interpretations of Deep Learning
A long line of work interprets neural networks through a Bayesian lens, from classical analyses of predictive uncertainty (MacKay, 1992; Neal, 2012) to variational or stochastic approximations of posterior inference (Graves, 2011; Blundell et al., 2015). Recent papers argue that, in large-data limits, minimizing cross-entropy implicitly targets the Bayesian posterior predictive (Xie et al., 2022; von Oswald et al., 2023). These results concern what training should produce at the population level. Our contribution is complementary: a controlled setting in which the true posterior is known, memorization is impossible, and one can directly test whether a finite transformer actually realizes this Bayesian computation.
@@ 7.2. In-Context Learning and Algorithmic Generalization
Transformers have been shown to perform algorithmic tasks in context, including arithmetic (Garg et al., 2022), synthetic induction (Elman, 1990), and more general pattern extrapolation (Akyürek and Andreas, 2022; Olsson et al., 2022). Behaviorally, these models often resemble Bayesian learners, an observation formalized by recent explanatory theories (Xie et al., 2022; von Oswald et al., 2023). However, prior work cannot distinguish true Bayesian computation from learned heuristics or memorized templates, because the ground-truth posterior is unknown for natural language tasks.
Our wind-tunnel methodology solves this identification problem: by constructing tasks with closed-form analytic posteriors and combinatorially large hypothesis spaces, we obtain a direct pointwise comparison between model predictions and Bayes’ rule. This moves the discussion from correlation to mechanism.

### v3

@@ 4.3. Length Generalization Requires Late-Layer Attention
To identify which components support stable rollout, we train a variant transformer in which attention is disabled in the top two layers but FFNs and residuals remain intact. The no-late-attention model fits the training horizon reasonably well ($1.57\times 10^{-3}$ bits), but breaks down under rollout:
@@
4.4. Associative Recall: Content-Based Retrieval
The associative recall task tests whether architectures can retrieve stored information by content—the binding primitive. Unlike bijection or HMM tasks where the relevant context is determined by position or dynamics, here the model must identify which cue–target pair to retrieve based on a probe that arrives only at test time. The transformer achieves 100% accuracy on held-out sequences, demonstrating perfect content-based retrieval. This requires the attention mechanism: the probe cue must match against stored cues to route information from the corresponding target. MLPs did not succeed on this task under our training protocol (accuracy at chance), as they process each position independently without cross-token interaction.
4.5. Continuous Bayesian Regression
[…]
4.6. Architectural Comparison: Which Primitives Does Each Architecture Realize?
To understand which architectural ingredients enable Bayesian inference, we decompose it into three inference primitives and test which architectures realize each:
  • Belief accumulation: integrating evidence into a running posterior (tested by bijection elimination)
  • Belief transport: propagating beliefs through stochastic dynamics (tested by HMM filtering)
  • Random-access binding: retrieving stored hypotheses by content (tested by associative recall)
Table 2 summarizes results across all three tasks.
Table 2. Architecture comparison across inference primitives. Bijection/HMM: entropy MAE in bits (lower is better; mean $\pm$ std over 5 seeds for Mamba/LSTM). Recall: accuracy (higher is better). Each task tests a different primitive. Transformers realize all three; Mamba realizes accumulation and transport; LSTMs realize only accumulation (of static sufficient statistics); MLPs realize none.
Architecture Bijection HMM Assoc. Recall Primitives
(accumulation) (transport) (binding) realized
Transformer 0.007 0.049 100% All 3
[…]
MLP 1.85 0.40 0 of 3
Transformers realize all three primitives.
Transformers achieve near-exact Bayesian posteriors on bijection and HMM, and perfect accuracy on associative recall with 64 cue–target pairs. Attention provides all three capabilities: accumulation via the residual stream, transport via content-based routing, and binding via query-key matching.
Mamba realizes accumulation and transport, but struggles with binding.
[…]
LSTMs realize only accumulation of static sufficient statistics.
The LSTM achieves 0.009 bits on bijection—comparable to Transformer and Mamba—but fails on both HMM (0.416 bits) and associative recall (0.5%, random chance). The key distinction: bijection elimination admits a static sufficient statistic (the set of observed outputs, representable in fixed dimension) that LSTM can track with its recurrent state. But HMM requires a sufficient statistic that evolves under dynamics—the belief vector must be transported through the transition matrix at each step—and associative recall requires the statistic to be indexed by content. Under our training protocol, LSTM’s fixed gating accumulated evidence but did not learn to implement the content-dependent operations required for transport or binding.
MLPs realize no primitives.
Without sequence structure, MLPs collapse to marginal predictions and fail uniformly.
The primitives taxonomy predicts performance.
This decomposition explains the full pattern of results: each architecture succeeds precisely on tasks requiring only the primitives it can implement. The taxonomy also resolves apparent contradictions in the literature—Mamba beats transformers on tasks dominated by transport (HMM) but loses on tasks requiring binding (recall).
Figure 2. Mamba discovers the 5-corner geometry of HMM belief space. Final-layer representations from Mamba on the HMM task (5 hidden states). Left: Points colored by most likely hidden state reveal five distinct clusters—one per state. Right: The same points colored by posterior entropy show confidence variation within each cluster (red = low entropy/high confidence, blue = high entropy/uncertainty). Mamba has learned that the belief simplex has five corners and organized its representations accordingly. This geometric structure explains its strong performance (MAE = 0.024 bits, 5-seed average).
@@ 4.7. Multi-Seed Consistency
To ensure that Bayesian tracking is not an artifact of initialization or optimization noise, we repeated transformer HMM experiments across five independent random seeds. Per-position error curves for all seeds (Figure 13) nearly overlap at $K=20$, $K=30$, and $K=50$. The seed-to-seed variability is negligible compared to the gap between architectures, confirming that the learned Bayesian algorithm is robust to initialization and training noise. The Mamba and LSTM HMM results in Table 2 are 5-seed averages with standard deviations of 0.009 and 0.003 bits respectively—confirming that the large performance gaps between architectures (e.g., LSTM’s 0.411 vs Mamba’s 0.024 bits on HMM) reflect genuine architectural differences rather than seed-to-seed variation.
Figure 3. Bijection wind tunnel: transformer matches the Bayesian posterior; MLP does not. Entropy trajectories at 150k training steps. The transformer lies essentially on top of the analytic Bayes curve across positions, while the capacity-matched MLP barely reduces uncertainty and fails to implement hypothesis elimination. This is the comparison summarized quantitatively in Table 3 and discussed in Section 4.1.
Figure 4. Bijection wind tunnel: per-sequence entropy dynamics. Eight randomly chosen bijections from the test set. Each panel shows transformer entropy (solid) and analytic Bayes entropy (dashed) as a function of position. The sawtooth pattern—discrete drops when mappings are revealed and collapses to (near) zero when previously seen inputs reappear—confirms that the transformer is performing stepwise hypothesis elimination, not merely matching the Bayes curve in aggregate.
@@
Figure 13. Multi-seed robustness of HMM length generalization. Overlay of per-position transformer MAE curves across five random seeds for $K=20$, $K=30$, and $K=50$. Seed-to-seed variability is negligible relative to the transformer–MLP gap, showing that the learned Bayesian algorithm is robust to initialization and optimization noise.
@@
Figure 18. Value-manifold unfurling during training. PCA projection of attention outputs in the bijection model, colored by analytic posterior entropy. At 100k steps, low-entropy states are tightly clustered; by 150k, they lie along a smooth one-dimensional curve parameterized by entropy, enabling fine-grained encoding of posterior states.Each point is an attention output (head output or block attention output – whichever you used) at a supervised prediction position; PCA is fit on the pooled outputs and then plotted, colored by analytic posterior entropy.
@@ 5.1. Layer 0 Creates the Hypothesis Frame
The computation begins with a structural operation: Layer 0 attention constructs the hypothesis space in which all subsequent inference takes place. Keys at this layer form an approximately orthogonal basis over input tokens (Figure 16), providing a coordinate system over which posterior mass can be represented and manipulated. We measure orthogonality via mean absolute off-diagonal cosine similarity between key vectors. Across 5 seeds, the bijection model achieves $0.052\pm 0.004$ versus $0.082\pm 0.003$ for random vectors in $d=192$ dimensions—a 37% reduction ($p<0.001$, paired $t$-test). The HMM model shows similar structure: $0.061\pm 0.006$ vs. $0.079\pm 0.002$ random baseline. Head-wise ablations confirm the indispensability of this step. A single Layer 0 “hypothesis-frame head” dominates the layer’s contribution (Figure 6), and ablating this head alone severely disrupts calibration. Here “hypothesis-frame head” means the head whose keys span the near-orthogonal basis over hypothesis tokens and whose values instantiate the corresponding per-hypothesis slots in the residual stream. No other attention head exhibits comparable sensitivity. This identifies a structural bottleneck: forming the hypothesis frame is a prerequisite for any later Bayesian computation. Once established, this frame remains stable through training. Attention maps at Layer 0 change little across checkpoints, even as the value manifold and calibration improve substantially. The model therefore learns the geometry of the inference problem early, and subsequently refines numerical precision within this fixed frame.
@@ 5.2. Sequential Bayesian Elimination Across Depth
Layer-wise ablations (Figure 5) show that removing any single layer (attention + FFN, as implemented) increases calibration error by more than an order of magnitude. This demonstrates that the computation is not shallow or redundant. Each layer provides a distinct and non-interchangeable refinement step, forming a sequential, compositional realization of Bayesian updates. Together, these observations indicate that transformers implement Bayesian elimination not via a single transformation, but through a depth-wise sequence of projections and refinements within the Layer 0 frame.
@@ 5.3. Attention as Content-Addressable Routing
Across all depths, attention serves a consistent geometric role: it retrieves the components of the belief state relevant for the next update. Three observations support this routing interpretation:
@@
  • Stable routing during late refinement (Figures 18 and 19) shows that once the frame is correct, attention maps change minimally even as calibration improves.
@@ 5.4. Value-Space Manifolds and Precision Refinement
After routing stabilizes, the final layers refine the precision of the posterior representation. Figures 18 and 19 show that:
@@
This geometric unfurling enables fine-grained encoding of posterior confidence and accounts for late-position improvements in calibration. This refinement occurs while attention maps remain nearly unchanged, producing a clear frame–precision dissociation: attention defines where information flows, while downstream transformations refine how precisely beliefs are encoded.
@@ 5.5. Synthesis: A Three-Stage Architectural Mechanism
5.6. Mamba’s Mechanism: Selective State-Space Dynamics
While the three-stage mechanism above is specific to transformer attention, Mamba achieves comparable HMM performance (0.024 vs 0.049 bits MAE) through qualitatively different mechanisms. Analyzing Mamba’s internal representations reveals how selective state-space dynamics implement belief transport without attention.
Five-cluster geometry emerges from state selection.
Mamba’s final-layer representations organize into five discrete clusters corresponding to the five HMM hidden states (Figure 2), with within-cluster variation encoding posterior entropy. This is the same corner geometry of the belief simplex that transformers learn, achieved through input-dependent state selection rather than query-key matching.
Distributed representations enable belief transport.
Layer-wise analysis (Figure 21) reveals a key difference between Mamba and LSTM:
  • Mamba: Final-layer representations achieve $R^{2}=0.40$ for entropy prediction, with PC1 explaining only 21.9% of variance—the model maintains a distributed, multi-dimensional belief representation.
  • LSTM: Final-layer representations achieve $R^{2}=0.004$ (near random), with PC1 explaining 92.3% of variance—the model collapses to a one-dimensional manifold uncorrelated with true posterior entropy.
This explains the primitives difference: under our training protocol, LSTM’s fixed gating did not maintain the multi-dimensional structure required for belief transport, while Mamba’s selection mechanism preserves it.
Layer-wise entropy encoding.
Unlike transformers where entropy correlation increases sharply in later layers, Mamba shows moderate correlation ($|r|\approx 0.11$) across middle layers with gradual improvement in linear predictability ($R^{2}$) from 0.025 at Layer 1 to 0.40 at Layer 9. This suggests Mamba accumulates information more gradually through its recurrent state rather than in discrete elimination steps.
Convergence to Bayesian geometry.
The fact that different mechanisms—attention’s query-key matching and Mamba’s selective state updates—converge to similar geometric solutions (corner geometry of the belief simplex) suggests that Bayesian geometry may be a universal attractor for architectures with content-based routing. This provides a geometric explanation for why both architectures succeed at belief transport despite using fundamentally different computational primitives.
Figure 21. Layer-wise analysis: Mamba vs LSTM on HMM. For each layer, we measure PC1 variance (dimensionality), entropy correlation (alignment with uncertainty), and linear $R^{2}$ (predictability of entropy). Left: LSTM collapses to a 1D manifold (PC1 >> 90%) with near-zero entropy correlation. Right: Mamba maintains distributed representations (PC1 $\approx$ 22%) with increasing entropy predictability across depth. This explains why LSTM fails at belief transport while Mamba succeeds.
@@ 6.2. Depth as Compositional Necessity
A central conclusion from the ablation studies is that depth is not redundant. In both wind tunnels, removing any individual layer increases calibration error by more than an order of magnitude (Figure 5). This shows that Bayesian reasoning is expressed as a sequence of compositional projections, each layer refining the belief state in a way that cannot be collapsed into a single transformation. This stands in contrast to wide, shallow architectures: even with comparable parameter counts and identical training, MLPs fail to perform hypothesis elimination or state tracking (Section 4.6). Bayesian inference requires hierarchical refinement, and transformers supply the appropriate inductive bias through depth and residual composition.
@@ 6.3. From Wind Tunnels to Natural Language
While the wind tunnels are deliberately simplified, they capture the essential structure of probabilistic inference: integrating evidence over time to update latent beliefs. Large language models operate in a far more complex setting, with high-dimensional latent spaces and ambiguous, multi-modal evidence. Yet the geometric ingredients observed here—orthogonal hypothesis axes, depth-wise refinement, and stable routing—are structural rather than task-specific. The results therefore suggest that the probabilistic behaviors exhibited by LLMs may arise not only from scale or data richness but also from architectural geometry. Wind tunnels provide a verifiable lower bound: they show that transformers can implement Bayesian inference exactly when the posterior is known.
6.4. The Primitives Taxonomy Explains Architectural Differences
The four-architecture comparison (Table 2) reveals that Bayesian inference is not monolithic—different tasks demand different inference primitives, and different architectures realize different subsets.
Why Mamba beats Transformer on HMM but loses on recall.
[…]
Why LSTM succeeds on bijection but fails everywhere else.
LSTM’s gates depend only on $(h_{t-1},x_{t})$—under our training protocol, they did not learn to perform content-based matching across positions. This suffices for accumulating static sufficient statistics (bijection admits a fixed-dimensional statistic: the set of observed outputs), but fails when:
  • The statistic must evolve under dynamics (transport): LSTM compresses HMM representations to a 1D manifold uncorrelated with the true 5D belief state—it did not transport the belief vector through the transition matrix
  • The statistic must be indexed by content (binding): LSTM achieves 0.5% on recall (random chance)—it did not retrieve by content
Implications.
The primitives framework provides a principled basis for architecture selection: match the architecture’s capabilities to the task’s primitive requirements. It also explains why attention remains essential for tasks requiring flexible retrieval, even as alternatives like Mamba excel at sequence modeling.
@@ 6.5. A Lower Bound for Reasoning in LLMs
The wind tunnels establish a principled baseline for mechanistic reasoning in transformers. If a model cannot implement Bayes in a setting with a closed-form posterior and impossible memorization, it offers little evidence of genuine inference capability in natural language. Conversely, the fact that small, verifiable transformers succeed here—with interpretable geometric mechanisms—suggests that similar structures may underpin reasoning in large models. This provides a concrete research direction: search for the same geometric signatures in frontier LLMs. The diagnostics used here—key orthogonality, QK sharpening, value-manifold structure, and routing stability—offer testable predictions for analyzing pretrained language models.
@@ 7.1. Bayesian Interpretations of Deep Learning
A long line of work interprets neural networks through a Bayesian lens, from classical analyses of predictive uncertainty (MacKay, 1992; Neal, 2012) to variational or stochastic approximations of posterior inference (Graves, 2011; Blundell et al., 2015). Recent papers argue that, in large-data limits, minimizing cross-entropy implicitly targets the Bayesian posterior predictive (Xie et al., 2022; von Oswald et al., 2023). These results concern what training should produce at the population level. Our contribution is complementary: a controlled setting in which the true posterior is known, memorization is computationally infeasible, and one can directly test whether a finite transformer actually realizes this Bayesian computation.
@@ 7.2. In-Context Learning and Algorithmic Generalization
Transformers have been shown to perform algorithmic tasks in context, including arithmetic (Garg et al., 2022), synthetic induction (Elman, 1990), and more general pattern extrapolation (Akyürek and Andreas, 2022; Olsson et al., 2022). Behaviorally, these models often resemble Bayesian learners, an observation formalized by recent explanatory theories (Xie et al., 2022; von Oswald et al., 2023). Müller et al. (2024) provide evidence that high-capacity transformers often mimic the Bayesian predictor during in-context learning, and Reuter et al. (2025) demonstrate that transformers can perform full Bayesian inference for statistical models including generalized linear models and latent factor models, achieving results comparable to expensive exact methods. However, prior work cannot distinguish true Bayesian computation from learned heuristics or memorized templates, because the ground-truth posterior is unknown for natural language tasks. Furthermore, these studies focus on whether transformers can implement Bayesian inference, not why they succeed where other architectures fail. Our wind-tunnel methodology addresses both gaps: by constructing tasks with closed-form analytic posteriors and combinatorially large hypothesis spaces, we obtain a direct pointwise comparison between model predictions and Bayes’ rule. By comparing multiple architectures, we identify the inference primitives that explain success and failure. This moves the discussion from correlation to mechanism, and from existence to architectural characterization.